OpenStudy (anonymous):

\[\int \frac{x^4}{(x-1)(x^2+1)}dx\]

OpenStudy (dumbcow):

Divide it out, then use partial fractions \[\rightarrow \int\limits_{}^{} x +1+\frac{1}{2(x-1)} -\frac{x+1}{2(x^{2}+1)} dx\]

OpenStudy (ash2326):

we have \[\int\limits_{}^{} (x^4)/(x-1)(x^2+1) dx\] here degree of numerator is greater than denominator so we'll transform into the form p(x)+q(x)/r(x) \[\int\limits_{}^{} (x^2(x^2-1)+x^2)/((x-1)(x^2+1)) dx\] \[\int\limits_{}^{} (x^2/(x-1)- (x^2+1-1)/((x-1)(x^2+1)\] \[\int\limits_{}^{} x^2/(x-1)-1/(x-1)+1/((x-1)(x^2+1)) dx\] combining the first two terms \[\int\limits_{}^{} (x+1)+ 1/((x-1)(x^2+1)) dx\] split the second term using partial fractions \[1/(x-1)(x^2+1)= (-x-1)/2(x^2+1)+ 1/2(x-1)\] so the integral is \[\int\limits_{}^{} (x+1)+(-x-1)/2(x^2+1) -1/2(x-1) dx\] \[\int\limits_{}^{} (x+1)-x/2(x^2+1)-1/2(x^2+1)-1/2(x-1) dx\] now integrating \[(x+1)^2/2-1/4 \ln(x^2+1)-1/2\tan^{-1} x-1/2 \ln(x-1) + c\] c= constant of integration

OpenStudy (dumbcow):

\[=\frac{1}{2}x^{2}+x+\frac{1}{2}\ln(x−1)+\frac{1}{4}\ln(x2+1)+\frac{1}{2}\tan^{−1}x+C\]

OpenStudy (mr.math):

This was too easy Ishaan, you should post another one! ;)